### v2

# Alleviating Non-identifiability: a High-fidelity Calibration Objective for Financial Market Simulation with Multivariate Time Series Data

Abstract
The non-identifiability issue has been frequently reported in the social simulation works, where different parameters of an agent-based simulation model yield indistinguishable simulated time series data under certain discrepancy metrics. This issue largely undermines the simulation fidelity yet lacks dedicated investigations. This paper theoretically analyzes that incorporating multiple time series data features in the model calibration phase can alleviate the non-identifiability exponentially with the increasing number of features. To implement this theoretical finding, a maximization-based aggregation function is applied to existing discrepancy metrics to form a new calibration objective function. For verification, the financial market simulation, a typical and complex social simulation task, is considered. Empirical studies on both synthetic and real market data witness the significant improvements in alleviating the non-identifiability with much higher simulation fidelity of the chosen agent-based simulation model. Importantly, as a model-agnostic method, it achieves the first successful simulation of the high-frequency market at seconds level. Hence, this work is expected to provide not only a rigorous understanding of non-identifiability in social simulation, but a high-fidelity calibration objective function for financial market simulations.
1 Introduction
Social simulation plays a pivotal role of understanding the complex social systems in a way of revealing their endogenous complexity  [1]. The core is to construct a parameterized simulation model that can generate the simulated time series data resembling the observed states over time of the social system. By properly modeling the endogenous components of the social system as multiple interacting agents, i.e., the Agent-based Modeling (ABM) techniques  [2], various “what-if” tests can be conducted  [3]. That is, by intervening certain agents of the simulation model, the corresponding endogeneity of the real social system is revealed by observing how the evolution of the simulation model is influenced by the intervention.
The success of a social simulation largely lies with the model fidelity. Note that, the what-if tests are often counterfactual, i.e., they do not happen in real social systems and thus there are no ground-truth for validation. To what extent should we believe the results of what-if tests? In social systems, due to the uncertainty of human activities, it is often impractical to derive trustworthy simulation models mathematically or physically. Usually, this issue is addressed in a data-driven manner via model calibration  [4].
The basic idea is that, if a simulation model $M(\omega)$ can generate simulated data $\mathbf{X}_{T}$ resembling the observed data $\hat{\mathbf{X}}_{T}$ of the targeted real system within the discrete time interval $T=[t_{s},t_{s}+1,\cdots,t_{e}]$, it is considered that the underlying probabilistic distribution of the real system for generating $\hat{\mathbf{X}}_{T}$ is learned by $M(\omega)$. Here, $\hat{\mathbf{X}}_{T}$ denotes the time series $\{\hat{\mathbf{X}}_{t}\}^{t=t_{e}}_{t_{s}}$ for short. Hence, while using $M(\omega)$ for what-if test, by keeping the non-intervened agents of $M(\omega)$ unchanged, the simulation results of how the intervened agent influences the system are considered trustworthy. To obtain such $M(\omega)$, the discrepancy between $\mathbf{X}_{T}$ and $\hat{\mathbf{X}}_{T}$ is measured by various metrics $D$  [5]. The smaller $D$ is, the closer $M(\omega)$ approximates the distribution of data generation behind the real social system. By fixing the ABM-based model structure $M$, calibrating the parameter $\omega$ to fit a given observed data sequence $\hat{\mathbf{X}}_{T}$ is often modeled as an optimization problem  [5]:
$\min_{\omega\in\Omega}D(\hat{\mathbf{X}}_{T},M(\omega)=\mathbf{X}_{T})$ (1)
In practice, the above data-driven calibration methods often suffer from the so-called non-identifiability issue  [6, 7]. That is, there are multiple parameters in the parameter space $\Omega$ enjoy the same discrepancy value, given the same $\hat{\mathbf{X}}_{T}$. However, those parameters highly likely lead to different distributions of data generation, due to the nonlinear model structure $M(\omega)$. In other words, the “ground-truth” parameter, deciding the “ground-truth” data generation distribution, cannot be effectively identified from the other parameters in those sets. Consequently, the what-if tests on the resultant model of a randomly picked non-identifiable parameter therefrom are less trustworthy.
Despite that the non-identifiability is frequently mentioned in the literature, there is neither a formal definition of it nor a rigorous way to alleviate it. This paper first defines the non-identifiability issue in the form of multivariate time-series data. It is then discussed that the non-identifiability arises from that the previously calibrated time series data $\hat{\mathbf{X}}_{T}$ is often univariate and is insufficient to describe the uniqueness of a social system. For simplicity and applicability, this work restricts the scenarios to the financial market simulation, as it is a representative social simulation and has been active for over 30 years  [8]. Like other complex social systems, the real-world financial markets will output multivariate states data over time, e.g., price, volume, bid/ask directions, order arrival time, and other hand-crafted features  [9]. However, existing works merely consider the observed price data for calibration, resulting in wide criticism for their low simulation fidelity  [10], especially in intraday simulations  [11, 6].
This paper alleviates the non-identifiability issue by incorporating more data features into the calibration process, each of which represents a univariate time series data. We mathematically show that the non-identifiability can be reduced exponentially with the number of features. To this end, the multivariate time series is fed into existing discrepancy metrics in a way of multiple univariate time series. That is, the data associated with each feature is calibrated individually. This results in multiple individual calibration tasks with respect to the multiple features, respectively. Note that, the non-identifiable set of the multivariate time series calibration task should be non-identifiable across all the individual tasks. Specifically, only the ground-truth parameter locates simultaneously at the global optimum of all parameter spaces of those individual tasks. To implement it, we propose to search in the intersection of all those individual spaces and implement it via a maximization function aggregating all those individual calibration tasks. Such an implementation enjoys a high scalability that one can incorporate as many as possible data features to alleviate the non-identifiability.
Extensive empirical studies have been conducted to show that the proposed method can significantly alleviate the non-identifiability issue of financial market simulation on 10 synthetic data and 1 real data from Shenzhen Stock Exchange of China. Such advantages are verified robust over 6 commonly seen financial market data features. Besides, this paper successfully reports the high-fidelity simulation results of the high-frequency market at seconds level, which is the first time for the ABM-agnostic calibration methods.
The rest of this paper is organized as follows. Section  2 describes the background of financial market simulation and related works. Section  3 theoretically discuss the alleviation of non-identifiability. Section  4 reports the empirical studies in details. Section  5 concludes this work.
2 Background
2.1 ABM-based Financial Market Simulation
In social simulations, Agent-based Modeling (ABM) based models are dominantly preferred than the black-box deep learning models due to their interpretability of how the data is generated  [12]. ABM constructs a collection of agents mimicking various social activities and interacting together to simulate the whole social system  [12]. In financial market simulation, the agents are designed to model various trading strategies of human  [13] (illustrated in Figure  1), e.g., fundamentalists, momentum traders, high-frequency traders, market makers, and arbitrage traders. Recent financial market simulation works favor the Preis-Golke-Paul-Schneid (PGPS) model, which consists of two general types of agents, i.e., the liquidity providers and the liquidity takers  [14].
Figure 1: An illustration of the ABM-based financial market simulator with various agents interacting with the simulated exchange.
Regardless of the differences among agents, the ABM-based financial market simulations follow the same workflow. That is, at the beginning of the simulation, the parameterized agents are initialized with some parameters, which determine how exactly those agents submit orders under their own simulated strategies. Each submitted order usually consists of the price, volume, direction, and arrival time. The exchange is simulated as a Limited Order Book (LOB) that maintains the existing untraded orders submitted by the agents  [15].
Figure 2: A LOB organizes all the untraded orders in terms of directions, prices, arrival time and volume.
Those untraded orders are categorized into two directions (bid or ask) and sorted by their prices. For the orders with the same price, they are again sorted by the arrival time. Each agent continuously observes the current LOB and submit its order at will. If a new order matches a price in the opposite direction of the LOB, it will be traded immediately, and the matched orders will be removed from the LOB  [16]. In existing financial market simulations, the mid-price (see Figure  2) is mostly considered as the observable time series data for calibration, as it implies the potential movement of the trading prices and is important to many what-if tests, e.g., the causes of flash crashes  [17, 18].
2.2 Calibration of Financial Market Simulation
To ensure the fidelity of a financial market simulation, the model often needs careful calibration of the agents’ parameters, so that the simulated data resembles the observed time series from the real market. The general form of calibration has been given in Eq. (1).
In recent years, the study of the discrepancy metrics $D$ has attracted increasing research attentions, as the distance measure between time series remains unsolved and impacts significantly on the calibration performance. Methods of Simulated Moments  [19] are a typical method that first transforms the time series into several statistical moments, and then calculates the weighted average between the moments of observed data and simulated data. The information criterion based objective functions are also studied  [20, 21]. These methods abstract the temporal information of the original time series via various techniques like histograms and measure the distances between the abstracted temporal vectors. The non-parametric Kolmogorov-Smirnov test is also introduced as $D$ by measuring the distance between the probabilistic distributions of the observed data and the simulated data  [7]. Others employ the tool of Bayesian theory to estimate the likelihood  [22] or posterior distribution  [23] of the observed data.
Despite of the progresses of exploiting various metrics for calibrating the financial market simulation, existing works frequently encounter the so-called non-identifiability issue, resulting in poor simulation fidelity  [24]. From our viewpoint, one major reason is that existing works merely use the mid-price data to calibrate the model, ignoring that the real financial market continuously output the data with multiple features. Although the mid-price might be important to the market participants, the other features describe the market in more details. Unfortunately, the above mentioned calibration objective functions are all designed for univariate time series data (for mid-price). Consequently, existing works are ill-equipped to calibrate with multivariate time series data. Yet, it is still unclear how exactly it benefits from calibrating more features.
3 A High-fidelity Calibration Objective
This section first formally defines the non-identifiability, then theoretically analyzes how it benefits from calibrating with multivariate data, and lastly proposes a new calibration objective function using multiple features.
3.1 The Definition of Non-identifiability
As existing discrepancy metrics $D$ are designed for univariate data, we re-denote the calibrated data as $\hat{\mathbf{X}}_{T}^{k}$ with a superscript $k$ to emphasize that it is a 1-dimensional time series describing some $k$-th feature of the real system. In financial markets, this feature can be either price, volume, bid/ask directions, order arrival time, or any other hand-crafted features  [9]. Correspondingly, the simulated data generated by $M(\omega)$ is re-denoted as $\mathbf{X}^{k}_{T}$. Then the original calibration task can be re-written as $\min_{\omega}D(\hat{\mathbf{X}}_{T}^{k},M(\omega)=\mathbf{X}^{k}_{T})$, and the non-identifiability issue is defined as follows.
Let $\Omega^{D,k}$ be the parameter space of this calibration task. There exists some region $\mathcal{S}^{D,k}\subseteq\Omega^{D,k}$ that any two different parameters $\omega_{1},\omega_{2}\in\mathcal{S}^{D,k}$ satisfy $D(\hat{\mathbf{X}}_{T}^{k},M(\omega_{1}))=D(\hat{\mathbf{X}}_{T}^{k},M(\omega_{2}))$. Such region $\mathcal{S}^{D,k}$ is called the non-identifiability set and is non-empty if the ground-truth parameter $\omega^{*}$ is assumed to be included in $\Omega^{D,k}$. Then the non-identifiability is defined as the probability of uniformly randomly sampling a parameter $\omega$ from $\Omega^{D,k}$ while falling into $\mathcal{S}^{D,k}$, denoted as $P(\omega\in\mathcal{S}^{D,k}|\hat{\mathbf{X}}_{T}^{k})$. Let $|\mathcal{S}|$ denote the cardinality of a set $\mathcal{S}$, it gives
$P(\omega\in\mathcal{S}^{D,k}|\hat{\mathbf{X}}_{T}^{k})=\frac{|\mathcal{S}^{D,k}|}{|\Omega^{D,k}|}$ (2)
Therefore, the rank of any $\omega_{1},\omega_{2}\in\mathcal{S}^{D,k}$ cannot be effectively identified by the objective $D$ given the data $\hat{\mathbf{X}}_{T}^{k}$, leading to a flat landscape of the parameter space therein and the search-based calibration process easily gets stuck in $\mathcal{S}^{D,k}$. Note that, $\mathcal{S}^{D,k}$ can involve arbitrary parameters in $\Omega^{D,k}$, including the “ground-truth” parameter $\omega^{*}$. Hence, if the calibration process converges to $\mathcal{S}^{D,k}$, one has to randomly pick one $\omega\in\mathcal{S}^{D,k}$ for further what-if simulation. Unfortunately, the underlying data generation distribution decided by any $\omega\in\mathcal{S}^{D,k}$ is less likely to be the same with $\omega^{*}$, given the nonlinear nature of the simulation model $M(\omega)$. Consequently, the what-if test on the randomly picked $\omega$ is less trustworthy.
3.2 Alleviating the Non-identifiability with Multiple Features
In this section, it is shown that the non-identifiability issue can be alleviated exponentially with $K>1$ features of the observed time series data.
Suppose the observed time series generated by the real social system is multivariate containing $K$ features, i.e., $\hat{\mathbf{X}}_{T}=\{\hat{\mathbf{X}}_{T}^{k}\}_{k=1}^{K}$, which is true for the financial markets. And the simulation model $M(\omega)$ can also generate a multivariate simulated data $\mathbf{X}_{T}=\{\mathbf{X}_{T}^{k}\}_{k=1}^{K}$ with the same $K$ features. Given that existing discrepancy metrics $D$ only computes for two univariate time series, there are in total $K$ individual calibration tasks: $\min_{w}D(\hat{\mathbf{X}}_{T}^{k},M(w)=\mathbf{X}_{T}^{k}),k=1,\cdots,K.$
Note that, given any two parameters $\omega_{1},\omega_{2}\in\Omega^{D,k}$, only they are simultaneously non-identifiable in all $K$ individual calibration tasks, i.e., $\omega_{1},\omega_{2}\in\{\mathcal{S}^{D,k}\}_{k=1}^{K}$, are they non-identifiable with the general calibration task with respect to $\hat{\mathbf{X}}_{T}$. Otherwise, their discrepancy to $\hat{\mathbf{X}}_{T}$ can be distinguishable in at least one feature and thus are identifiable. That is to say, only the intersection of $\{\mathcal{S}^{D,k}\}_{k=1}^{K}$ can be regarded as the non-identifiable set in the multivariate time series setting. Hence, the non-identifiability in the multivariate time series setting can be calculated as:
$P(\omega\in\{\mathcal{S}^{D,k}\}_{k=1}^{K}|\hat{\mathbf{X}}_{T})=\frac{|\cap_{k=1}^{K}\mathcal{S}^{D,k}|}{|\Omega^{D}|}$ (3)
It is easy to obtain that $\frac{|\cap_{k=1}^{K}\mathcal{S}^{D,k}|}{|\Omega^{D}|}\leq\frac{|\mathcal{S}^{D,k}|}{|\Omega^{D,k}|},k=1,..,K$. Hence, we have:
$P(\omega\in\{\mathcal{S}^{D,k}\}_{k=1}^{K}|\hat{\mathbf{X}}_{T})\leq P(\omega\in\mathcal{S}^{D,k}|\hat{\mathbf{X}}_{T}^{k})$ (4)
In Eq.(3) and Eq.(4), the equality holds only if the $K$ non-identifiable sets are fully overlapped. This can be caused by two reasons. First, the $K$ features are fully dependent that the time series data of one feature can be deduced from other features. This suggests that the selected features should be as diverse as possible. Second, the parameter spaces of all the other $K$ individual tasks are full of non-identifiable sets, i.e., $\mathcal{S}^{D,k}=\Omega^{D,k},k=1,2,...,K$, which is highly unlikely with existing discrepancy metrics $D$. Therefore, with proper choice of the features and $D$, the non-identifiability of calibrating the multivariate time series $\hat{\mathbf{X}}_{T}$ is smaller than that of the univariate $\hat{\mathbf{X}}_{T}^{k}$. In other words, calibrating multiple features can alleviate the non-identifiability issue. Next, we estimate how fast it decreases with $K$ as follows.
For clarity, let $A_{k}$ denote the event of a uniformly randomly sampled parameter falling into the non-identifiable set of the $k$-th individual calibration task. The probability of happening $A_{k}$ is $P(A_{k})=P(\omega\in\mathcal{S}^{D,k}|\hat{\mathbf{X}}_{T}^{k})$. Then the non-identifiability in the multivariate time series setting can be rewritten as:
$P(\omega\in\{\mathcal{S}^{D,k}\}_{k=1}^{K}|\hat{\mathbf{X}}_{T})$
$=$ $P(A_{1},A_{2},...,A_{K})$
$=$ $P(A_{1})P(A_{2}|A_{1})P(A_{3}|A_{2},A_{1})\cdots P(A_{K}|A_{K-1},...,A_{1})$
[…]
3.3 Aggregating Multiple Features via Maximization
Note that, the above discussions require the multivariate time series $\{\hat{\mathbf{X}}_{T}^{k}\}_{k=1}^{K}$ being calibrated separately as $K$ univariate time series with existing discrepancy metrics $D$. This results in $K$ discrepancy values of $D(\hat{\mathbf{X}}_{T}^{k},M(w)=\mathbf{X}_{T}^{k}),k=1,\cdots,K$. This section discusses how to optimally aggregate the $K$ values, so that the practical alleviation of the non-identifiability follows the above analysis.
Let $F$ denote the proposed objective function for calibrating the multivariate $\hat{\mathbf{X}}_{T}$. To achieve $\mathcal{S}^{F}=\cap_{k=1}^{K}\mathcal{S}^{D,k}$, the proposed way is to ensure every location in $\Omega^{F}$ to be the worst among all $K$ parameter space $\{\Omega^{D,k}\}_{k=1}^{K}$ (see Figure  3 for illustration). In this case, any parameter $\omega$ that is not simultaneously in all $K$ non-identifiable sets will make it fall outside of the $\mathcal{S}^{F}$. This is because that the parameters outside the non-identifiable set is worst than those who are non-identifiable from the ground-truth parameter $\omega^{*}$.
Given that the calibration problem is a minimization problem, the new objective function $F$ can be implemented with a maximization aggregation among $K$ individual calibration tasks with $D$. That is,
$\min_{\omega\in\Omega}F(\hat{\mathbf{X}}_{T},M(w)=\mathbf{X}_{T})$ (6)
$=\min_{\omega\in\Omega}$ $\max_{k\in\{1,\cdots,K\}}D(\hat{\mathbf{X}}_{T}^{k},M(w)=\mathbf{X}_{T}^{k})$
Figure 3: Optimizing the maximization of two functions is equivalent to the search in the intersection between the two search spaces.
4 Empirical Studies
This section empirically verifies two major claims of this paper.
  • With the increasing number of features applied in the calibration, the non-identifiable set shrinks largely.
  • The fidelity of the simulation data, in terms of various performance metrics, is improved significantly by calibrating more features.
4.1 Experimental Settings
To conduct a calibration process, 4 terms should be specified according to Eq. (6), i.e., the discrepancy metric $D$, the simulation model $M(\omega)$, the optimization algorithm $\rm{min}_{\omega}$, and the $K$ features $\{\hat{\mathbf{X}}_{T}^{k}\}_{k=1}^{K}$.
The discrepancy metric. The Wasserstein distance is employed as $D$ since it provides geometric interpretability, and tail sensitivity  [25, 26]. Given two univariate time series $\mathbf{X}_{T}^{k}$ and $\hat{\mathbf{X}}_{T}^{k}$, the Wasserstein distance is defined as
$D(\hat{\mathbf{X}}_{T}^{k},\mathbf{X}_{T}^{k})=\int_{-\infty}^{+\infty}|U(x)-V(x)|\rm{d}\textit{x}$ (7)
where
$U(x)=\frac{1}{T}\sum_{t=t_{s}}^{t_{e}}\mathcal{I}(\hat{\mathbf{X}}_{t}^{k}\leq x)$ (8)
$V(x)=\frac{1}{T}\sum_{t=t_{s}}^{t_{e}}\mathcal{I}(\mathbf{X}_{t}^{k}\leq x)$
Here $U$ and $V$ are the empirical cumulative distribution functions for the data set $\hat{\mathbf{X}}_{T}^{k}$ and $\mathbf{X}_{T}^{k}$, respectively. And $\mathcal{I}(\cdot)$ is an indicator function that returns 1 if the input event is true, otherwise it returns 0.
The simulation model. This work considers the Preis-Golke-Paul-Schneid (PGPS) model as the simulation model, due to its popularity in recent financial market simulation works  [11, 6]. The PGPS model simulates the market dynamics through interactions between two types of agents, i.e., 125 liquidity providers and 125 liquidity takers. At the $t$-th time-step, each liquidity provider submits a limited order at a fixed probability $\alpha$ with the default volume equals 1. The probability of a limited order being bid side or ask side equals to 0.5. Each liquidity taker submits a market order at a fixed probability $\mu$ with the default volume equals 1. The probability of a market order being bid side or ask side is $q_{taker}(t)$ or $1-q_{taker}(t)$, respectively. The probability $q_{taker}(t)$ is specified by a mean-reverting random walk with mean equals 0.5, mean reversion probability equals $0.5+|q_{taker}(t)-0.5|$, and the increment sizes towards mean equals ±$\pm\Delta_{s}$. Moreover, the liquidity taker has a probability $\delta$ to cancel its untraded limited order. Let $p_{a}(t)$ and $p_{b}(t)$ represents the best ask and bid price, respectively. The price of a market order is determined by the market, i.e., the best price at the opposite side of the LOB. The price of a limited order is determined as $p_{a}(t)+(\lambda(t)\log(u))+s$, where $s=1$ if it is a ask order and otherwise $s=-1$, and $u\sim U(0,1)$. Here $\lambda(t)$ is a time-variant order placement depth parameter and is calculated as
$\lambda(t)=\lambda_{0}(1+\frac{|q_{taker}(t)-\frac{1}{2}|}{\sqrt{<q_{taker}(t)-\frac{1}{2}>^{2}}}C_{\lambda})$ (9)
$\sqrt{<q_{taker}(t)-\frac{1}{2}>^{2}}$ is a pre-computed value. Before initiating the simulation, it is computed in advance through $10^{5}$ Monte Carlo iterations, ensuring its convergence to the accurate standard deviation.
To summarize, the PGPS model has 6 hyper-parameters shared by all agents that need to be calibrated, i.e., $\omega=[\delta,\lambda_{0},C_{\lambda},\Delta_{s},\alpha,\mu]$.
The optimization method. Since the order matching process needs to align with the real market, the PGPS model should be run on the executable programs based some trading simulators [^2] This work builds the simulation environment based on the Multi-Agent Exchange Environment (MAXE) simulator developed by University of Oxford  [27].. As a result, the calibration objective function is non-differentiable and the calibration process often has to resort to non-derivative optimization methods. In such contexts, both Bayesian optimization methods and evolutionary algorithms are particularly suitable. This work employs the well-known Particle Swarm Optimization (PSO) algorithm  [28] to optimize $\omega$ with respect to Eq.(6). For simplicity, the standard version of PSO  [29] is considered. Without further fine-tuning, the hyper-parameters of PSO follow the suggested configurations, where the population size is set to 40, the inertia weight is set to 0.8, the cognitive and social crossover parameters $c_{1}=0.5,c_{2}=0.5$.
The selected features. Six different yet commonly seen features are considered  [15, 30], which mainly concern either the price or volume information of the LOB data and are listed as follows:
  • $f_{1}$, the mid-price $m_{t}=\frac{p_{a}(t)+p_{b}(t)}{2}$ at each $t$-th step;
  • $f_{2}$, the total traded volume within each $t$-th step;
  • $f_{3}$, the mid-price return within each $t$-th step $\ln m_{t+1}-\ln m_{t}$;
  • $f_{4}$, the spread at each $t$-th step $p_{a}(t)-p_{b}(t)$;
  • $f_{5}$, the total volume of the best bid price at each $t$-th step;
  • $f_{6}$, the total volume of the best ask price at each $t$-th step.
By incorporating the 6 selected features, 6 objective functions are constructed based on Eq.(6), denoted as $F_{i},i=1,2,...,6$. Each $F_{i}$ indicates that the first $i$ features are used for calibration, i.e., $F_{i}=\max_{k\in\{1,\cdots,i\}}D(\hat{\mathbf{X}}_{T}^{k},M(\omega))$. In the following empirical studies, it is shown how the increasing number of features influences the simulation fidelity.
Parameters $\mathbf{\delta}$ $\mathbf{\lambda_{0}}$ $\mathbf{C_{\lambda}}$ $\mathbf{\Delta_{S}}$ $\mathbf{\alpha}$ $\mathbf{\mu}$
data 1 0.025 100 10 0.001 0.15 0.025
data 2 0.002 200 10 0.002 0.1 0.03
data 3 0.05 152 45 0.003 0.13 0.05
data 4 0.02 288 27 0.003 0.11 0.04
data 5 0.04 62 35 0.003 0.12 0.04
data 6 0.01 171 15 0.0015 0.15 0.03
data 7 0.033 114 14 0.0033 0.1 0.047
data 8 0.01 129 30 0.0017 0.05 0.03
data 9 0.02 62 2 0.003 0.14 0.02
data 10 0.01 87 23 0.001 0.09 0.05
Table 1: Parameter settings of generating 10 synthetic data with the PGPS model.
The test protocol. The test data contains 11 targeted time series to be calibrated, including 10 synthetic data and 1 real data. The 10 synthetic time series are generated by setting 10 different parameters of the PGPS model as shown in Table  1, each of which consists of 3600 time steps at the frequency of 1 second. The real data is chosen as 000001.sz from Shenzhen Stock Exchange of China, consisting of 1200 time steps at a frequency of 3 seconds from 9:30 a.m. to 10:30 a.m. of a day in 2019. The time budget of $10^{4}$ simulation evaluations is allowed for the calibration of each targeted data. That is, the PSO runs for 250 iterations with population size 40, and the best parameter with the lowest objective function value is applied into PGPS model to simulate data as $\mathbf{X}_{T}$. The performance of the calibration is quantitatively validated using the Wasserstein (W) distance and the Mean Square Error (MSE) commonly seen in time series analysis. Both metrics are applied to measure the discrepancy between the synthetic data and simulated data on each feature, normalized with the min-max method, respectively.
(a) $F_{1}$
(b) $F_{2}$
(c) $F_{3}$
(d) $F_{4}$
(e) $F_{5}$
(f) $F_{6}$
Figure 4: With the increasing number of features, the cardinality of the non-identifiable set decreases significantly.
4.2 Results and Analysis
Effectiveness of multiple features. To show how the increasing numbers of features alleviate the non-identifiability issue, we depict the changes of the 2-dimensional parameter space of $[\alpha,\mu]$. Specifically, a target data of 3600 time steps is generated using the PGPS model with the commonly recommended default parameter setting $\omega^{*}=[0.025,100,10,0.001,0.15,0.025]$  [6]. On this default setting, 10000 combinations are created by sampling in $[\alpha,\mu]$ in grid. By calculating the objective value $F$ between the target data and each of the 10000 combinations, it obtains the objective values of the 10000 combinations in the space of $[\alpha,\mu]$, which are depicted as Figure  4.(a)-(f). The 6 figures in Figure  4.(a)-(f) are depicted according to the objective values of $F_{1}$, $F_{2}$,…,$F_{6}$, respectively. Only the combinations whose objective values rank in the best 10% over all the corresponding features, are they depicted in blue, otherwise they will be depicted in yellow. For example, in Figure  4.(c), the combinations depicted in blue are those whose discrepancy values rank in 10% among the 10000 combinations on all the first 3 features $\{f_{i}\}_{i=1}^{3}$. By this means, the blue points indicate the intersection of the combinations who rank best 10% in all the corresponding features in $F$. In this regard, the non-identifiable set are also depicted in blue as they are indistinguishable with $\omega^{*}$.
As can be seen that, with the increaseha of the features, the blue set, as well as the non-identifiable set, decreases significantly. When 5 to 6 features are included in the objective function, there leaves very few blue points, indicating that the non-identifiability issue is indeed alleviated very effectively.
Note that, although the non-identifiable set of any $F_{i}$ is verified to be smaller than that of any $F_{j}$, $i>j$, it does not deterministically lead to better calibration performance of $F_{i}$ than $F_{j}$. Because $F$ mainly decides the parameter space, while the calibration performance is also jointly influenced by the optimization algorithms. Given the calibration problem is black-box with multiple local optima, it is non-trivial for the optimization algorithms to find the global optimum even the parameter space is modeled fully identifiable.
data $obj$ $F_{1}$ $F_{2}$ $F_{3}$ $F_{4}$ $F_{5}$ $F_{6}$
Data 1 W 7.49 3.17 4.85 6.25 2.59 4.90
MSE 6.97 4.46 3.93 6.25 4.38 4.04
Data 2 W 3.68 1.58 2.60 5.81 4.40 3.88
MSE 4.15 4.18 4.88 5.25 5.15 4.33
Data 3 W 2.34 5.52 6.74 4.79 5.95 8.07
MSE 5.03 4.00 4.42 3.62 3.58 4.38
Data 4 W 1.17 1.99 1.72 4.54 1.98 5.70
MSE 3.26 3.04 3.65 3.19 3.36 3.90
Data 5 W 3.97 3.66 3.03 7.88 7.65 2.54
MSE 4.85 3.56 3.65 6.08 4.01 3.66
Data 6 W 6.15 2.52 7.34 5.98 2.42 7.45
MSE 5.65 4.4 5.62 5.87 4.54 6.44
Data 7 W 5.66 5.10 3.29 2.38 0.99 1.99
MSE 5.20 3.76 3.88 3.71 3.55 3.72
Data 8 W 2.64 0.88 2.32 3.88 3.96 2.23
MSE 3.11 2.78 3.48 3.90 4.12 2.81
Data 9 W 9.28 1.88 3.10 8.74 8.71 4.42
MSE 4.08 3.86 7.43 7.65 5.02 3.71
Data 10 W 1.25 5.40 2.12 0.99 0.57 0.90
MSE 2.88 3.68 2.58 2.64 2.65 2.62
#win W 0 7 6 5 5 6
MSE 0 8 6 5 6 7
Table 2: Calibrating with the multivariate time series ($F_{2}$-$F_{6}$) robustly outperforms that with the univariate data ($F_{1}$).
Calibration of the synthetic data. For each of the 10 synthetic data, by optimizing the 6 objective functions $F_{i},i=1,...,6$ with PSO, it results in 6 calibrated models and 6 simulated data. Note that, even though a model is calibrated using $F_{i},i<6$, it can still functionally generate the simulated data $\mathbf{X}_{T}$ with respect to all 6 features.
Figure 5: The simulated data using the proposed method resembles the synthetic data on all 6 features.
All 6 discrepancy values $D$ between each simulated data and its corresponding synthetic data is measured by both the W and MSE, based on the 6 features, respectively. In Table  2, each cell displays the average of these 6 discrepancy values. It shows that by using more than one feature ($F_{2}$ to $F_{6}$) for calibration, it consistently outperforms the calibration with the univariate data ($F_{1}$). Specifically, on the synthetic data of data 1, 3, and 7, calibrating with $F_{1}$ is defeated by all of the other 5 objective functions who use at least two features. On other instances, there are at least two objectives among $F_{2}$ to $F_{6}$ outperform $F_{1}$.
To a further step, the numbers of synthetic data that each $F_{i},i=2,...,6$ outperforms $F_{1}$ are counted and recorded in the last row of Table  2. It clearly shows that each $F_{i},i=2,...,6$ wins $F_{1}$ on at least half of the testing instances on both W and MSE metrics. More specifically, $F_{2}$ wins the most among $F_{2}$ to $F_{6}$. Note that, all these 5 objectives employ the second feature of total traded volume. It can be attributed that this feature contributes the most to the calibration of multiple features. This is reasonable as the traded volume is a data feature very different to the first feature of mid-price.
Figure 6: By calibrating with more features, the simulated data resembles the real market data better.
Then, we depict the high-frequency time series of 6 features to show how the simulated data resembles the synthetic data in Figure  5. Specifically, the simulated data calibrated by $F_{5}$ on data 7 is displayed as it represents the best calibration performance that $F_{2}$ to $F_{6}$ all outperform $F_{1}$ on both the Wasserstein distance and MSE metric. It can be seen that the simulated time series resemble the synthetic data on all 6 features very closely. To our best knowledge, this is the first ABM-agnostic method of successfully calibrating the high-frequency market time series at 1 second level and simulating 6 features with high-fidelity. In comparison, existing works mostly calibrate the mid-price data at a daily frequency [6].
Calibration of the real data. At last, it is verified that the proposed method can also calibrate the high-frequency real market data at a 3-second level. A 1-hour time series of 000001.sz is calibrated using $F_{1}$, $F_{2}$, and $F_{3}$. The simulated data generated by the corresponding best calibrated models is depicted in Figure  6, together with the real observed data. Due to the limited space, only the mid-price time series is shown in a time step of 1 minute. It can be seen that the simulated mid-price of $F_{1}$ is less similar to the real observed data. By incorporating more features, the fidelity of the simulated mid-price improves by $F_{2}$ and $F_{3}$. This confirms that the proposed method indeed can calibrate the real market data with multiple time series, thus suggesting a new effective calibration objective function for financial market simulation.
5 Conclusions
This paper studies the non-identifiability issue of social simulation. To our best knowledge, this is the first time that the non-identifiability is formally defined and analyzed to be alleviated. It is defined as the probability of randomly sampling in the non-identifiable set of the parameter space. It is rigorously analyzed that this probability can be reduced with more time series features in an exponential manner. Based on the analysis, a new objective function is proposed to effectively aggregate multiple features with a maximization function over existing discrepancy metrics. Extensive empirical studies have been conducted on 10 synthetic data and 1 real data of the financial market simulation tasks. It has been successfully verified in the simulations that, with multiple features, not only the non-identifiable set can be minimized but the simulation fidelity is largely improved. Hence, this paper provides a high-fidelity objective function to calibrate the high-frequency financial market simulator.