### v3

# Alleviating Non-identifiability: a High-fidelity Calibration Objective for Financial Market Simulation with Multivariate Time Series Data

Abstract
The non-identifiability issue has been frequently reported in the social simulation works, where different parameters of an agent-based simulation model yield indistinguishable simulated time series data under certain discrepancy metrics. This issue largely undermines the simulation fidelity yet lacks dedicated investigations. This paper theoretically analyzes that incorporating multiple time series data features in the model calibration phase can alleviate the non-identifiability exponentially with the increasing number of features. To implement this theoretical finding, a maximization-based aggregation function is proposed based on existing discrepancy metrics to form a new calibration objective function. For verification, the task of calibrating the Financial Market Simulation (FMS), a typical yet complex social simulation, is considered. Empirical studies confirm the significant improvements in alleviating the non-identifiability of calibration tasks. Furthermore, as a model-agnostic method, it achieves much higher simulation fidelity of the chosen FMS model on both synthetic and real market data. Hence, this work is expected to provide not only a rigorous understanding of non-identifiability in social simulation, but an off-the-shelf high-fidelity calibration objective function for FMS.
1 Introduction
Social simulation plays a pivotal role of understanding the complex social systems in a way of revealing their endogenous complexity  [27]. By properly modeling the endogenous components of the social system as multiple interacting agents, i.e., using the Agent-based Models (ABMs)  [10], various “what-if” tests can be conducted by intervening the agents of interest and analyzing the evolution process of the simulators  [8, 29].
The success of a social simulation largely lies with the model fidelity. Note that, the what-if tests are often counterfactual, i.e., they do not happen in real social systems and thus there are no ground-truth for validation. To what extent should we believe the results of what-if tests? Besides, due to the uncertainty of human activities, it is often impractical to derive mathematically trustworthy social simulation models. Usually, this issue is addressed in a data-driven manner via model calibration  [16].
Specifically, a simulation model $M(\omega)$ is essentially a data generative process, i.e., $M(\omega)=\mathbf{X}_{T^{\prime}},T^{\prime}=[1,...,\tau],\tau\in\mathbb{N}^{+}$. If the simulated data $\mathbf{X}_{T}$ resembles the observed data $\hat{\mathbf{X}}_{T}$ of the targeted real system within any discrete time interval of interest $T=[t_{s},t_{s}+1,...,t_{e}]$, it is considered that the underlying data generating probabilistic distribution of the real system for generating $\hat{\mathbf{X}}_{T}=\{\hat{\mathbf{X}}_{t}\}^{t_{e}}_{t=t_{s}}$ is learned by $M(\omega)$. Hence, while using $M(\omega)$ for what-if test, by keeping the non-intervened agents of $M(\omega)$ unchanged, the simulation results of how the intervened agent influences the system are considered trustworthy. To obtain such $M(\omega)$, the discrepancy between $\mathbf{X}_{T}$ and $\hat{\mathbf{X}}_{T}$ is measured by various metrics $D$  [24]. The smaller $D$ is, the closer $M(\omega)$ approximates the data generating distribution of the real social system. By fixing the ABM-based model structure $M$, calibrating the parameter $\omega$ to fit a given observed data sequence $\hat{\mathbf{X}}_{T}$ is often modeled as an optimization problem  [24]:
$\min_{\omega\in\Omega^{D}}D(\hat{\mathbf{X}}_{T},M(\omega)=\mathbf{X}_{T})$ (1)
In practice, the calibration methods often suffer from the so-called non-identifiability issue  [11, 1]. That is, there are multiple parameters in the parameter space $\Omega^{D}$ share the same discrepancy value to the given $\hat{\mathbf{X}}_{T}$. However, those parameters highly likely lead to different distributions of data generation, due to the nonlinear model structure $M(\omega)$. In other words, the “ground-truth” parameter of deciding the “ground-truth” data generating distribution cannot be effectively identified from the other parameters in those sets. Consequently, the what-if tests on the resultant model of a randomly picked non-identifiable parameter therefrom are less trustworthy.
Despite that the non-identifiability is frequently mentioned in the literature, there is neither a formal definition to describe it nor a rigorous way to alleviate it. This paper first defines the non-identifiability as the probability of a randomly uniformly sampled parameter falling into the non-identifiable set. With this definition, this paper mathematically proves that the non-identifiability can be alleviated by calibrating with more distinct features, each of which represents an observed univariate time series. More specifically, we mathematically show that the upper bound of the non-identifiability can be reduced exponentially with the number of features, if the selected features satisfy certain conditions.
This explains well that the previous works mostly calibrate their models to merely one data feature and is thus less sufficient to describe the uniqueness of a social system. For clarification and applicability, this work restricts the scenarios to the Financial Market Simulation (FMS), as it is a representative social simulation and has been active for over 30 years  [22]. Like other complex social systems, the real-world financial markets will output multivariate states data over time, e.g., price, volume, bid/ask directions, order arrival time, and other hand-crafted features  [21]. However, existing works merely consider the observed mid-price data for calibration, resulting in wide criticism for their low simulation fidelity  [9], especially in the high-frequency intraday simulations  [25, 11].
Note that, following our theory, the task of calibrating to multivariate time series is naturally divided into multiple individual tasks, each of which calibrates the model to one distinct univariate time series. In practice, how to jointly execute those individual tasks so that the obtained parameter is more identifiable to the original calibration tasks with multivariate time series? This challenges the implementation of the above theory. To address it, we propose a novel calibration objective by aggregating all those individual calibration tasks using the maximization function. This new objective function mathematically asks to search in the intersection of all those individual spaces, thus the obtained parameter can be highly identifiable across all individual tasks and thus fit our theory accurately.
Extensive empirical studies have successfully shown that the proposed method can significantly alleviate the non-identifiability issue of financial market simulation. Such advantages are verified robust over 6 commonly seen financial market data features. Besides, this paper also presents the high-fidelity simulation results by using the proposed new calibration objective function on 10 synthetic data and 1 real data from Shenzhen Stock Exchange of China.
The reminder of this paper is as follows. Section 2 describes the background of financial market simulation and related works. Section 3 theoretically discuss the alleviation of non-identifiability. Section 4 reports the empirical studies in details. Section 5 concludes this work.
2 Background
2.1 Calibration of Social Simulators
To ensure the simulation fidelity, the simulator often needs careful calibration of the agent parameters, so that the simulated data resembles the observed time series from the real system. The general form of calibration refers to Eq. (1).
In recent years, the study of the discrepancy metrics $D$ has attracted increasing research attentions, as the distance measure between time series remains unsolved and impacts significantly on the calibration performance. Methods of Simulated Moments  [12] are a typical method that first transforms the time series into several statistical moments, and then calculates the weighted average between the moments of observed data and simulated data. The information criterion based objective functions  [18, 2] abstract the temporal information of the original time series via various techniques like histograms and measure the distances between the abstracted temporal vectors. The non-parametric Kolmogorov-Smirnov test is also introduced as $D$ by statistically testing between the probabilistic distributions of the observed data and the simulated data  [1]. Other methods employ the Bayesian theory to estimate the likelihood  [17] or posterior distribution  [13] of the observed data also require a well-defined $D$ for selecting effective samples and updating the estimated distributions.
Figure 1: An illustration of the ABM-based FMS.
Despite of the progresses of exploiting various metrics for calibrating social simulators, existing works frequently encounter the so-called non-identifiability issue, resulting in poor simulation fidelity  [19]. From our viewpoint, one major reason is that existing works merely use the one single feature to calibrate the model, ignoring that the real systems continuously output multivariate data. Although the used feature might be a major factor to describe the social system, the other features can help capture more details of the system. Unfortunately, existing calibration objectives are mostly designed for univariate time series data. Consequently, existing works are ill-equipped to calibrate with multivariate time series data. Yet, it is theoretically unclear how exactly it benefits from calibrating more features.
2.2 ABM-based FMS and Data Features
In social simulations, ABMs are dominantly preferred than the black-box deep learning models due to their interpretability of the data generating process  [14, 20]. ABM constructs a collection of agents mimicking various social activities and interacting together to simulate the whole social system  [23]. In FMS, the agents are designed to model various trading strategies  [4] (illustrated in Figure 1), e.g., fundamentalists, momentum traders, high-frequency traders, market makers, and arbitrage traders.
Regardless of the differences among agents, the ABM-based FMSs follow the same workflow. That is, at the beginning of the simulation, the parameterized agents are initialized with the calibrated parameters, which determine how exactly those agents submit orders under their own simulated strategies. Each submitted order usually consists of the price, volume, direction, and arrival time. The Limited Order Book (LOB), a special data structure for organizing the existing untraded orders, is maintained by the exchange and can be observed by all trading agents. The simulated exchange generates the next LOB by matchmaking the incoming orders submitted by the agents and the current LOB  [28].
Those untraded orders are categorized into two directions (bid or ask) and sorted by their prices. For the orders with the same price, they are again sorted by the arrival time. Each agent continuously observes the current LOB and submit its order at will. If a new order matches a price in the opposite direction of the LOB, it will be traded immediately, and the matched orders will be removed from the LOB. Undoubtedly, many data features can be extracted from LOB  [21]. However, existing FMSs mostly considered the mid-price as the only feature for calibration, which represents the average between the best ask price and best bid prices (see Figure 2) and implies the potential price movement of the market.
Figure 2: LOB contains much information about the market.
3 A High-fidelity Calibration Objective
This section first formally defines the non-identifiability, then theoretically analyzes how it benefits from calibrating with multivariate data, and lastly proposes a new calibration objective using multiple features to implement our theory.
3.1 The Definition of Non-identifiability
As existing discrepancy metrics $D$ are designed for univariate data, we re-denote the target data as $\hat{\mathbf{X}}_{T}^{k}$ with a superscript $k$ to emphasize that it is a 1-dimensional time series describing some $k$-th feature of the real system. In financial markets, this feature can be either price, volume, bid/ask directions, order arrival time, or any other hand-crafted features  [21]. Correspondingly, the $k$-th feature of the simulated data is re-denoted as $\mathbf{X}^{k}_{T}$. Let $\Omega^{D,k}$ be the parameter space of this calibration task. Then the original calibration task can be re-written as
$\min_{\omega\in\Omega^{D,k}}D(\hat{\mathbf{X}}_{T}^{k},M(\omega)=\mathbf{X}^{k}_{T}).$ (2)
Without loss of generality, it is assumed that the optimal parameter $\omega^{*}$ lies in the defined parameter space, i.e., $\omega^{*}\in\Omega^{D,k}$, where $D(\hat{\mathbf{X}}_{T}^{k},M(\omega^{*}))<\epsilon$. Note that, the optimal parameter may not be unique by definition. The cardinality of the optimal set is influenced by $\epsilon$, which is used to filter the observation noise of data. Then the non-identifiable set is defined as follow:
Definition 1 (The non-identifiable set)
There exists some region $\mathcal{S}^{D,k,\epsilon}\subseteq\Omega^{D,k}$ that for all $\omega_{1},\omega_{2}\in\mathcal{S}^{D,k,\epsilon}$ such that $D(\hat{\mathbf{X}}_{T}^{k},M(\omega_{1}))<\epsilon$ and $D(\hat{\mathbf{X}}_{T}^{k},M(\omega_{2}))<\epsilon$. Such region $\mathcal{S}^{D,k,\epsilon}$ is called the non-identifiable set.
The cardinality of $\mathcal{S}^{D,k,\epsilon}$, denoted as $|\mathcal{S}^{D,k,\epsilon}|$, is also influenced by $\epsilon$. Then the non-identifiability can be defined as follows:
Definition 2 (The univariate non-identifiability)
Let $P(\omega\in\mathcal{S}^{D,k,\epsilon}|\hat{\mathbf{X}}_{T}^{k})$ be the non-identifiability to the $k$-th univariate time series data. It it defined as the probability of uniformly randomly sampling a parameter $\omega$ from $\Omega^{D,k}$ while it falls into $\mathcal{S}^{D,k,\epsilon}$. And we have
$P\left(\omega\in\mathcal{S}^{D,k,\epsilon}\middle|\hat{\mathbf{X}}_{T}^{k}\right)=\frac{|\mathcal{S}^{D,k,\epsilon}|}{|\Omega^{D,k}|}.$ (3)
Therefore, the rank of any $\omega_{1},\omega_{2}\in\mathcal{S}^{D,k,\epsilon}$ cannot be effectively identified by the objective $D$ given the data $\hat{\mathbf{X}}_{T}^{k}$, leading to that the comparison-based calibration process easily gets stuck in $\mathcal{S}^{D,k,\epsilon}$. Note that, $\mathcal{S}^{D,k,\epsilon}$ can involve arbitrary parameters in $\Omega^{D,k}$, including the “ground-truth” parameter $\omega^{*}$. Hence, if the calibration process converges to $\mathcal{S}^{D,k,\epsilon}$, one has to randomly pick one $\omega\in\mathcal{S}^{D,k,\epsilon}$ for further what-if simulation. Unfortunately, the underlying data generative distribution decided by any randomly picked $\omega\in\mathcal{S}^{D,k,\epsilon}$ is less likely to be the same with $\omega^{*}$, given the nonlinear nature of the simulation model $M(\omega)$. Consequently, the what-if test on the randomly picked $\omega$ is less trustworthy, unless the cardinality of $\mathcal{S}^{D,k,\epsilon}$ is sufficiently small.
3.2 Alleviating the Non-identifiability with Multiple Features
In this section, it is shown that the non-identifiability issue can be alleviated exponentially with $K>1$ features of the observed time series data.
Suppose the observed time series generated by the real social system is multivariate containing $K$ features, i.e., $\hat{\mathbf{X}}_{T}=\{\hat{\mathbf{X}}_{T}^{k}\}_{k=1}^{K}$, which is true for the financial markets. And the simulation model $M(\omega)$ can also generate a multivariate simulated data $\mathbf{X}_{T}=\{\mathbf{X}_{T}^{k}\}_{k=1}^{K}$ with the same $K$ features. Given that existing discrepancy metrics $D$ only computes for two univariate time series, there are in total $K$ individual calibration tasks: $\min_{w}D(\hat{\mathbf{X}}_{T}^{k},M(w)=\mathbf{X}_{T}^{k}),k=1,\cdots,K.$
Note that, given any two parameters $\omega_{1},\omega_{2}\in\Omega^{D,k}$, only they are simultaneously non-identifiable in all $K$ individual calibration tasks, i.e., $\omega_{1},\omega_{2}\in\{\mathcal{S}^{D,k,\epsilon}\}_{k=1}^{K}$, are they non-identifiable in the general calibration task with respect to multivariate $\hat{\mathbf{X}}_{T}$. Otherwise, their discrepancy to $\hat{\mathbf{X}}_{T}$ can be distinguished in at least one feature and thus are identifiable. The search space $\Omega^{D,k}$ is identical across all calibration tasks and is uniformly represented as $\Omega^{D}$. That is to say, only the intersection of the $K$ individual non-identifiable set $\{\mathcal{S}^{D,k,\epsilon}\}_{k=1}^{K}$ can be regarded as the non-identifiable set in the multivariate time series setting, which is defined as
Definition 3 (The multivariate non-identifiability)
Let $P\left(\omega\in\{\mathcal{S}^{D,k,\epsilon}\}_{k=1}^{K}\middle|\hat{\mathbf{X}}_{T}\right)$ be the non-identifiability to the multivariate time series with $K$ features. It it defined as the probability of uniformly randomly sampling a parameter $\omega$ from $\Omega^{D}$ while it falls into all $K$ individual non-identifiable set $\{\mathcal{S}^{D,k,\epsilon}\}_{k=1}^{K}$. And we have
$P\left(\omega\in\{\mathcal{S}^{D,k,\epsilon}\}_{k=1}^{K}\middle|\hat{\mathbf{X}}_{T}\right)=\frac{\left|\cap_{k=1}^{K}\mathcal{S}^{D,k,\epsilon}\right|}{\left|\Omega^{D}\right|}.$ (4)
Since $\frac{\left|\cap_{k=1}^{K}\mathcal{S}^{D,k,\epsilon}\right|}{\left|\Omega^{D}\right|}\leq\min_{k}\frac{\left|\mathcal{S}^{D,k,\epsilon}\right|}{\left|\Omega^{D,k}\right|}$, we have $P\left(\omega\in\{\mathcal{S}^{D,k,\epsilon}\}_{k=1}^{K}\middle|\hat{\mathbf{X}}_{T}\right)\leq\min_{k}P\left(\omega\in\mathcal{S}^{D,k,\epsilon}\middle|\hat{\mathbf{X}}_{T}^{k}\right)$. This equality holds only if the $K$ non-identifiable sets are fully overlapped. This can be caused that the $K$ features are fully dependent that the time series data of one feature can be deduced from other features. This suggests that the selected features should be as diverse as possible. Therefore, with proper choice of the features and $D$, the non-identifiability of calibrating the multivariate time series $\hat{\mathbf{X}}_{T}$ is smaller than that of the univariate $\hat{\mathbf{X}}_{T}^{k}$. In other words, calibrating multiple features can alleviate the non-identifiability issue. Next, we estimate how fast it decreases with $K$ as follows.
For clarity, let $A_{k}$ denote the event of a uniformly randomly sampled parameter falling into $\mathcal{S}^{D,k,\epsilon}$. The probability of happening $A_{k}$ is $P(A_{k})=P(\omega\in\mathcal{S}^{D,k,\epsilon}|\hat{\mathbf{X}}_{T}^{k})$. Then we have:
$P\left(\omega\in\{\mathcal{S}^{D,k,\epsilon}\}_{k=1}^{K}\middle|\hat{\mathbf{X}}_{T}\right)$
$=P\left(A_{1}A_{2}\cdots A_{K}\right)$
$=P(A_{1})\cdot P(A_{2}|A_{1})\cdots P(A_{K}|A_{K-1}\cdots A_{1})$
$=\frac{\left|\mathcal{S}^{D,1,\epsilon}\right|}{\left|\Omega^{D}\right|}\cdot\frac{\left|\cap_{i=1}^{2}\mathcal{S}^{D,i,\epsilon}\right|}{\left|\mathcal{S}^{D,1,\epsilon}\right|}\cdots\frac{\left|\cap_{i=1}^{K}\mathcal{S}^{D,i,\epsilon}\right|}{\left|\cap_{i=1}^{K-1}\mathcal{S}^{D,i,\epsilon}\right|}$
[…]
Theorem 1 (The exponential alleviation)
If each $K$-th feature is selected such that $\beta_{K}^{2}<\frac{{\sum}_{i=1}^{K-1}\beta_{i}^{2}}{K-1}$, the upper bound of the multivariate non-identifiability is reduced exponentially with the increasing number of features $K$.
[…]
3.3 Aggregating Multiple Features via Maximization
Note that, the above discussions require the multivariate time series $\{\hat{\mathbf{X}}_{T}^{k}\}_{k=1}^{K}$ being calibrated separately as $K$ univariate time series with existing discrepancy metrics $D$. This results in $K$ individual calibration tasks of $D(\hat{\mathbf{X}}_{T}^{k},M(w)=\mathbf{X}_{T}^{k}),k=1,\cdots,K$. How to jointly calibrate the parameter to satisfy the $K$ individual tasks simultaneously? This section proposes a new objective function based on the $K$ individual objectives, so that the practical alleviation of the non-identifiability follows the above theory.
Let $F$ be the proposed new objective function used for calibrating the multivariate time series $\hat{\mathbf{X}}_{T}$, and let $\mathcal{S}^{F,\epsilon}$ denote the associated non-identifiable set defined by this objective function. To achieve $\mathcal{S}^{F,\epsilon}=\cap_{k=1}^{K}\mathcal{S}^{D,k,\epsilon}$, the proposed way is to ensure every location in $S^{F}$ to be the worst value among all $K$ non-identifiable sets $\{\mathcal{S}^{D,k,\epsilon}\}_{k=1}^{K}$ (see Figure 3 for illustration). In this case, any parameter $\omega$ that is not simultaneously in all $K$ non-identifiable sets will make it fall outside of the $\mathcal{S}^{F,\epsilon}$. This is because that the parameters outside the non-identifiable set is even worse than those who are non-identifiable from the ground-truth parameter $\omega^{*}$.
Figure 3: Optimizing the maximization of two functions equals to searching in their intersections.
Given that the calibration problem is a minimization problem, the new objective function $F$ can be implemented with a maximization aggregation among $K$ individual calibration tasks with $D$. Formally, we have the following theorem.
Theorem 2 (The utility of new objective function $F$)
Let $F$ aggregate the $K$ individual calibration tasks via maximization, i.e.
$F(\hat{\mathbf{X}}_{T},M(w)=\mathbf{X}_{T})$ (6)
$=$ $\max_{k\in\{1,\cdots,K\}}D(\hat{\mathbf{X}}_{T}^{k},M(w)=\mathbf{X}_{T}^{k}),$
then minimizing $F$ achieves $\mathcal{S}^{F,\epsilon}=\cap_{k=1}^{K}\mathcal{S}^{D,k,\epsilon}$.
The proof of Theorem 2 is provided in appendix.
4 Empirical Studies
This section empirically verifies two major claims.
  • With the increasing number of features applied in the calibration, the non-identifiable set shrinks exponentially.
  • The fidelity of the simulation data, in terms of various performance metrics, is improved significantly by calibrating more features.
4.1 Experimental Settings
To conduct a calibration process, 4 terms should be specified according to Eq. (6), i.e., the discrepancy metric $D$, the simulation model $M(\omega)$, the optimization algorithm $\rm{min}_{\omega}$, and the $K$ features $\{\hat{\mathbf{X}}_{T}^{k}\}_{k=1}^{K}$.
The discrepancy metric. The Wasserstein distance is employed as $D$ since it provides geometric interpretability, and tail sensitivity  [6, 5]. Given two univariate time series $\mathbf{X}_{T}^{k}$ and $\hat{\mathbf{X}}_{T}^{k}$, the Wasserstein distance is defined as
$D(\hat{\mathbf{X}}_{T}^{k},\mathbf{X}_{T}^{k})$ (7)
$=$ $\int_{-\infty}^{+\infty}\left|\frac{1}{T}\sum_{t=t_{s}}^{t_{e}}\mathcal{I}(\hat{\mathbf{X}}_{t}^{k}\leq x)-\frac{1}{T}\sum_{t=t_{s}}^{t_{e}}\mathcal{I}(\mathbf{X}_{t}^{k}\leq x)\right|\,\mathrm{d}x,$
where $\mathcal{I}(\cdot)$ is an indicator function that returns 1 if the input event is true, otherwise it returns 0.
The simulation model. This work considers the Preis-Golke-Paul-Schneid (PGPS) model as the simulation model, due to its popularity in recent financial market simulation works  [25, 11]. The PGPS model simulates the market dynamics through interactions between two types of agents, i.e., 125 liquidity providers and 125 liquidity takers. The PGPS model has 6 hyper-parameters shared by all agents that need to be calibrated, i.e., $\omega=[\delta,\lambda_{0},C_{\lambda},\Delta_{s},\alpha,\mu]$. The detailed descriptions of the behaviors of the agents and how the hyper-parameters control them can be found in Appendix.
The optimization method. Since the order matchmaking process needs to align with the real market, the PGPS model should be run on the executable trading programs. As a result, the calibration objective function is non-differentiable and the calibration process often has to resort to non-derivative optimization methods. In such contexts, both Bayesian optimization methods and evolutionary algorithms are particularly suitable. This work employs the well-known Particle Swarm Optimization (PSO) algorithm  [15] to optimize $\omega$ with respect to Eq. (6). The setting of PSO is in Appendix.
The selected features. Six different yet commonly seen features are considered  [28, 7], which mainly concern either the price or volume information of the LOB data and are listed as follows:
  • $f_{1}$, the mid-price $m_{t}=\frac{p_{a}(t)+p_{b}(t)}{2}$ at each $t$-th step;
  • $f_{2}$, the total traded volume within each $t$-th step;
  • $f_{3}$, the price return at each $t$-th step $\ln m_{t+1}-\ln m_{t}$;
  • $f_{4}$, the spread at each $t$-th step $p_{a}(t)-p_{b}(t)$;
  • $f_{5}$, the total volume of the best bid price at each $t$-th step;
  • $f_{6}$, the total volume of the best ask price at each $t$-th step.
By incorporating the 6 selected features, 6 objective functions are constructed based on Eq. (6), denoted as $F_{i},i=1,2,...,6$. Each $F_{i}$ indicates that the first $i$ features are used for calibration, i.e., $F_{i}=\max_{k\in\{1,\cdots,i\}}D(\hat{\mathbf{X}}_{T}^{k},M(\omega))$. In the following empirical studies, it is shown how the increasing number of features influences the simulation fidelity.
The test protocol. The test data contains 11 targeted time series to be calibrated, including 10 synthetic data and 1 real data. The 10 synthetic time series are generated by setting 10 different parameters of the PGPS model as shown in Appendix, each of which consists of 3600 time steps at the frequency of 1 second. The real data is chosen as 000001.sz from Shenzhen Stock Exchange of China, consisting of 1200 time steps at a frequency of 3 seconds from 9:30 a.m. to 10:30 a.m. of a day in 2019. The time budget of $10^{4}$ simulation evaluations is allowed for the calibration of each targeted data. That is, the PSO runs for 250 iterations with population size 40, and the best parameter with the lowest objective function value is applied into PGPS model to simulate data as $\mathbf{X}_{T}$. The performance of the calibration is quantitatively validated using the Wasserstein (W) distance and the Mean Square Error (MSE) commonly seen in time series analysis. Both metrics are applied to measure the averaged discrepancy between the synthetic data and simulated data on 6 features, i.e., $\frac{\sum_{k=1}^{6}D(\hat{\mathbf{X}}_{T}^{k},M(w)=\mathbf{X}_{T}^{k})}{6}$, where $D$ here indicates W or MSE. Each feature is normalized with the min-max method, respectively.
4.2 Results and Analysis
Figure 4: The illustration of the exponentially decreased non-identifiable set (grey dots). The red dot is the optima.
Alleviation of non-identifiability. Figure 4 illustrates the changes of the non-identifiable sets in the 2-dimensional parameter space of $[\alpha,\mu]$. This is done by three steps: 1) a target data of 3600 time steps is generated using the PGPS model with the recommended default parameter setting $\omega^{*}=[0.025,100,10,0.001,0.15,0.025]$  [11]; 2) On this default setting, 10000 combinations are created by sampling in $[\alpha,\mu]$ in grid while keeping the values of other 4 parameters unchanged; 3) By using $F$ to measure the descrepancy between the target data and the simulated data of each of the 10000 combinations, it obtains the objective values of the 10000 combinations on each $F_{i}$, which are depicted as Figure 4. The 6 figures in Figure 4 are depicted according to the objective values of $F_{1}$, $F_{2}$,…,$F_{6}$, respectively. Only the combinations whose objective values smaller than $\epsilon=0.1$ are depicted in grey, otherwise they will be depicted in blue. For example, in the third figure marked with $F_{3}$, the combinations depicted in grey are those whose discrepancy values smaller than $\epsilon=0.1$ on all the first 3 features $\{f_{i}\}_{i=1}^{3}$. By this means, the grey points indicate the non-identifiable set as they are indistinguishable with $\omega^{*}$ (marked as the red dot). As can be seen that, with the increase of the features, the grey non-identifiable set, decreases significantly. When 4 are incorporated in $F$, there exists only the optimum, indicating that the non-identifiability issue is indeed alleviated very effectively.
[…]
Figure 5: The non-identifiability (NI) decreases exponentially with the increasing features.
Figure 6: The distances of the simulated data to the target data decrease almost linearly with the increasing number of features incorporated in $F$.
Calibration to the synthetic data. Note that, although the non-identifiable set of any $F_{i}$ is verified to be smaller than that of any $F_{j}$, $i>j$, it does not deterministically lead to better calibration performance of $F_{i}$ than $F_{j}$. Because $F$ mainly decides the parameter space, while the calibration performance is also jointly influenced by the optimization algorithms. Given the calibration problem is black-box with multiple local optima, it is non-trivial for the optimization algorithms to find the global optimum even the parameter space is modeled fully identifiable.
For each of the 10 synthetic data, by optimizing the 6 objective functions $F_{i},i=1,...,6$ with PSO, it results in 6 calibrated models and 6 simulated data. Note that, even though a model is calibrated using $F_{i},i<6$, it can still functionally generate the simulated data $\mathbf{X}_{T}$ with respect to all 6 features.
The performance between each simulated data and its corresponding synthetic data is measured by the indicators based on both MSE and W, respectively. In Figure 6, each (blue or yellow) point displays the average performance on these 10 synthetic data. It shows that by using more features in $F$, the performance of calibration generally improves as the distance values decrease.
Furthermore, it is interesting to see how the simulation model calibrated by all features ($F_{6}$) compares to the ones calibrated to each single feature $f_{i},i=1,...,6$, instead of in an incremental manner of $F_{i},i=1,...,5$. As listed in Table 1, by calibrating with $F_{6}$, the obtained parameter enjoys the best simulation performance, i.e., the smallest averaged distances on 7 out of the 10 instances, and top 3 averaged distances on the other 3 instances. Besides, the friedman test is applied to show that $F_{6}$ achieves the best average performance and the advantage is statistically significant. This implies that the proposed theory and new objective function are applicable to various features and not restricted by the feature with more information of the social system, e.g., the mid-price.
$f_{1}$ $f_{2}$ $f_{3}$ $f_{4}$ $f_{5}$ $f_{6}$ $F_{6}$
Data 1 3.99 5.07 6.22 4.62 4.92 3.95 3.07
Data 2 3.65 3.95 4.67 3.73 6.18 4.77 2.90
Data 3 2.86 4.43 3.73 3.87 5.75 4.62 3.08
Data 4 2.83 3.50 3.47 3.87 4.33 2.99 2.47
Data 5 2.72 2.70 4.93 5.98 3.17 5.83 2.91
Data 6 5.23 5.58 4.02 6.67 4.48 6.77 3.67
Data 7 3.08 3.57 3.06 3.65 3.60 3.86 3.51
Data 8 4.88 4.50 2.56 3.82 3.91 3.62 2.26
Data 9 4.31 3.59 5.07 4.19 5.73 3.82 3.51
Data 10 2.53 2.43 2.95 6.93 3.22 2.24 2.23
#rank 3.20 3.60 3.80 4.50 5.30 3.80 1.57
Table 1: Comparisons of calibration performance between 6 univariate data and $F_{6}$. Note that the performance is at the $1E10^{-2}$ level.
Then, we illustrate the time series of 6 features to show how the simulated data resembles the synthetic data in Figure 7. Specifically, we calibrate the model by $F_{6}$ and $f_{x}$ on the data generated based on the recommended parameter setting $\omega^{*}=[0.025,100,10,0.001,0.15,0.025]$  [11], where $f_{x}$ indicates any single feature who helps to calibrate the model and generate the best simulated data. And the simulated data calibrated by $F_{6}$ resembles the target data on all 6 features very closely. To our best knowledge, this is the first ABM-agnostic method of successfully calibrating the high-frequency market time series at 1 second level and simulating 6 features with high-fidelity. In comparison, the simulated data of $f_{x}$ performs much poorer.
Figure 7: The simulated data using $F_{6}$ resembles the synthetic data on all 6 features.
Calibration to the real data. At last, it is verified that the proposed method can also calibrate the real market data at a 3-second level. A 1-hour time series of 000001.sz is calibrated using $F_{1}$ and $F_{3}$. The simulated data generated by the corresponding best calibrated models is depicted in Figure 8, together with the real observed data. Note that, most of the FMS simulators (including the adopted PGPS model) model the order volume in a coarse-grained manner, i,e., simply setting all the orders with the identical volume of 100. Hence, only the mid-price time series is shown in a time step of 1 minute. It can be seen that the simulated mid-price of $F_{3}$ is more similar to the real observed data than $F_{1}$ by an improvement of 27.5% in MSE. This confirms that the proposed method indeed can calibrate the real market data with multiple time series, thus suggesting a new effective calibration objective function for financial market simulation.
Figure 8: By calibrating with more features, the simulated data resembles the real market data better.
5 Conclusions
This paper studies the non-identifiability issue of social simulation. To our best knowledge, this is the first work that the non-identifiability is formally defined and theoretically alleviated. It is defined as the probability of randomly sampling in the non-identifiable set of the parameter space. It is rigorously analyzed that this probability can be reduced with more time series features in an exponential manner. Based on the analysis, a new objective function is proposed to effectively aggregate multiple features with a maximization function over existing discrepancy metrics. Extensive empirical studies have been conducted on 10 synthetic data and 1 real data of the FMS tasks. It has been successfully verified in the simulations that, with multiple features, not only the non-identifiable set can be minimized but the simulation fidelity is largely improved.